OpenStudy (anonymous):

What is a solution of x^2 +6x = -5

OpenStudy (anonymous):

factorize it @veritos

OpenStudy (anonymous):

just get all the terms on one side equating to zero..split the term 6x so that the sum is 6 and product is 5 and then simplify

OpenStudy (anonymous):

The answer can be -1?

OpenStudy (anonymous):

x^2 + 6x + 5 =0 factorize x^2 + 5x +x + 5 =0 x(x +5) 1(x + 5) =0 x+5 =0, x= -5 and x+1=0, x =-1

OpenStudy (anonymous):

Thank you very much!,,

OpenStudy (anonymous):

Can you help me with this solving problema please...